OpenStudy (anonymous):

how is d/dx e^-x^2 = -2xe^-x^2

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}e^{-x^2}=e^{-x^2}\frac{d}{dx}(-x^2)\]Chain rule! :)

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}e^{-x^2}=e^{-x^2}(-2x)\]Understand how that works?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you